OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the following system. 2x-3y=5 4x+2y=7

OpenStudy (anonymous):

do you know what is "x determinant" ?

OpenStudy (anonymous):

nopee :/

OpenStudy (anonymous):

I think you can take your 2x2 matrix = A. If that is true then you can use ad-bc to solve for the determinant. I'm not familiar with seeing 2x2 = something then figuring out determinant.

OpenStudy (anonymous):

im still confused..

OpenStudy (unklerhaukus):

\[|\mathbf M|=\begin{vmatrix}a&b\\c&d\end{vmatrix}=(a\times d)-(c\times b)\]

OpenStudy (tkhunny):

It is usually this: \(D = \left|\begin{matrix}2 & -3 \\ 4 & 2\end{matrix}\right|\) \(D_{x} = \left|\begin{matrix}5 & -3 \\ 7 & 2\end{matrix}\right|\) \(D_{y} = \left|\begin{matrix}2 & 5 \\ 4 & 7\end{matrix}\right|\) See where the constant terms are substituted for the coefficients in the subscripted versions.

OpenStudy (anonymous):

okay thank you